# Goal-guided Generative Prompt Injection Attack on Large Language Models

### v1

@@ Abstract
Current large language models (LLMs) provide a strong foundation for large-scale user-oriented natural language tasks. A large number of users can easily inject adversarial text or instructions through the user interface, thus causing LLMs model security challenges. Although there is currently a large amount of research on prompt injection attacks, most of these black-box attacks use heuristic strategies. It is unclear how these heuristic strategies relate to the success rate of attacks and thus effectively improve model robustness. To solve this problem, we redefine the goal of the attack: to maximize the KL divergence between the conditional probabilities of the clean text and the adversarial text. Furthermore, we prove that maximizing the KL divergence is equivalent to maximizing the Mahalanobis distance between the embedded representation $x$ and $x^{\prime}$ of the clean text and the adversarial text when the conditional probability is a Gaussian distribution and gives a quantitative relationship on $x$ and $x^{\prime}$. Then we designed a simple and effective goal-guided generative prompt injection strategy (G2PIA) to find an injection text that satisfies specific constraints to achieve the optimal attack effect approximately. It is particularly noteworthy that our attack method is a query-free black-box attack method with low computational cost. Experimental results on seven LLM models and four datasets show the effectiveness of our attack method.
Keywords:
Prompt Injection KL-divergence Robustness Mahalanobis Distance.
1 Introduction
@@
Large Language Models (LLMs) [19] are rapidly evolving in terms of architecture and functionality. As they become more deeply integrated into complex systems, the urgency to review their security properties also increases. Previous work has shown that even LLMs whose instructions are tuned through reinforcement learning with human feedback (RLHF) are vulnerable to adversarial attacks. Therefore, it is significant to study adversarial attacks on large language models. It can help researchers understand the security and robustness of large language models [26] to design more robust models to prevent such attacks.
@@
Previous attacks on large models mainly include white-box attacks and black-box attacks. White-box attacks assume that the attacker has full access to the model weights, architecture, and training workflow so that the attacker can obtain the gradient signal. The main method is gradient-based attacks. For example, Guo et al. [10] proposed a gradient-based distributed attack (GBDA), which, on the one hand, uses the Gumbel-Softmax approximation technique to make the adversarial loss optimization differentiable and, on the other hand, uses BERTScore and perplexity to enhance perceptibility and fluency. However, these methods can only attack open-source large language models but are powerless against more widely used closed-source LLMs such as ChatGPT.
@@
A black-box attack limits the attacker to access only API-type services. Black box attacks [7] can be divided into letter level, word level, sentence level and multi-layer level according to the attack granularity. Many black-box attacks use word replacement [12] to identify the most critical words in the text and replace them, or use some simple and general text enhancement methods [18], including synonym replacement, random insertion, random exchange or random deletion. Black-box strategies cannot know the internal structure of large models, so most attack methods use heuristic strategies. However, it is not clear how these heuristic strategies are related to the success rate of the attack, so they cannot effectively improve the success rate of the attack.
@@ 2 Related Work
In this section, we mainly review white-box and black-box attacks. Gradient-based attacks are in the white-box category, while jailbreak prompting, token manipulation, and prompt injection are all black-box attacks.
2.1 Gradient-based Attack
@@
Gradient-based Distributional Attack (GBDA) [10] uses the Gumbel-Softmax approximation technique to make the adversarial loss optimization differentiable while using BERTScore and perplexity to enhance perceptibility and fluency. HotFlip [4] maps text operations into vector space and measures the loss derivatives with respect to these vectors. AutoPrompt [27] uses a gradient-based search strategy to find the most effective prompt templates for different task sets. Autoregressive stochastic coordinate ascent (ARCA) [13] considers a broader optimization problem to find input-output pairs that match a specific pattern of behaviour. Wallace et al. [29] propose a gradient-guided search of markers to find short sequences, with 1 marker for classification and 4 additional markers for generation, called universal adversarial trigger (UAT), to trigger the model to produce a specific prediction. However, most of the widely used LLMs are not open source, so gradient-based white-box attacks are not suitable for these large language models. Our work belongs to the black-box attack method.
2.2 Jailbreak Prompting Attack
Jailbreak prompts hostile triggering of LLM to output harmful content. Jailbreaking is a black-box attack, so the wording combination is based on heuristics and manual exploration. Wei et al. [31] proposed two failure modes of LLM security to guide the design of jailbreak attacks: competing objectives and mismatched generalization. They also experiment with many jailbreak methods, including combined strategies such as the composition of prefix injection, style injection and the Base64 attack, etc. When observing prompt injection attacks, Greshak et al. [8] found that even if the attack does not provide detailed methods but only targets, the model may have access to more information that brings more risks such as phishing, private probing, and even proprietary information. It is worth noting that the goal of a jailbreaking prompting attack is to guide LLM to generate answers that exceed the safe range or alignment, while our attack method is a prompt injection attack, and the goal is to insert an adversarial prompt to make LLM generate wrong answers.
2.3 Token Manipulation Attack
@@
Given a text input containing a sequence of tokens, we can apply simple word operations (such as replacing them with synonyms) to trigger the model to make incorrect predictions. Ribeiro, et al. [24] manually define heuristic Semantic Equivalent Adversary Rules (SEAR) to perform minimal labelling operations, thereby preventing the model from generating correct answers. In contrast, Easy Data Augmentation (EDA) [32] defines a set of simple and more general operations to enhance text including synonym replacement, random insertion, random exchange, or random deletion. Given that context-aware prediction is a natural use case for masked language models, BERT-Attack [15] replaces words with semantically similar words via BERT. Unlike token manipulation attacks, our attack attempts to insert a generated adversarial prompt rather than just modifying certain words in the prompt.
2.4 Prompt Injection Attack
@@
Larger LLMs have superior instruction following capabilities and are more susceptible to these types of operations, which makes it easier for an attacker [17] to embed instructions in the data and trick the model into understanding it. Perez&Ribeiro [21] divides the targets of prompt injection attacks into goal hijacking and prompt leaking. The former attempts to redirect LLM’s original target to a new target desired by the attacker, while the latter works by convincing LLM to expose the application’s initial system prompt. However, system prompts are highly valuable to companies because they can significantly influence model behaviour and change the user experience. Liu et al. [16] found that LLM exhibits high sensitivity to escape and delimiter characters, which appear to convey an instruction to start a new range within the prompt. Our generative prompt injection attack method does not attempt to insert a manually specified attack instruction but attempts to influence the output of LLM by generating a confusing prompt based on the original prompt.
2.5 Black-box Attack Paradigm
@@
The black box attack paradigm refers to a broad range of attack methods that exploit vulnerabilities in machine learning models. Boundary query free attack BadNets [9] and boundary query related attack [6] are two common black box attack methods, whose difference lies in whether the boundaries of the model need to be input in advance to observe the output of the model. Other black box attack methods include methods based on model substitution such as practical black box attack[20] and context injection-based attacks, e.g. BERT-Attack [15].
3 Methodology
3.1 Threat Model with Black-box Attack
3.1.1 Adversarial scope and goal.
@@ 3.1 Threat Model with Black-box Attack
Given a text $t$ containing multiple sentences, we generate a text $t^{\prime}$ to attack the language model, ensuring that the meaning of the original text $t$ is preserved. Here, we use $\mathcal{D}(t^{\prime},t)$ to represent the distance between the semantics of text $t$ and $t^{\prime}$. If the LLM outputs $M(t)$ and $M(t^{\prime})$ differ, then $t^{\prime}$ is identified as an adversarial input for $M$. Our objective is to generate a text $t^{\prime}$ that satisfies the following conditions
$M(t)=r,\quad M(t^{\prime})=r^{\prime},\quad r\neq r^{\prime},\quad\mathcal{D}(t^{\prime},t)=0,$ (1)
where the texts $r$ and $r^{\prime}$ are the outputs of model $M$ on text $t$ and $t^{\prime}$ respectively, and $r$ is also the groundtruth of text $t$. Note that here we assume that the output text $r$ and $r^{\prime}$ come from some candidate set, so equality $=$ makes sense. In our problem, we aim to provide the following attack characteristics
@@
  • Effective: Problem 1 shows that the attack model ensures a high attack success rate (ASR) with $M(t^{\prime})=r^{\prime}$ such that $r^{\prime}\neq r$ on one hand and maintains high benign accuracy with $M(t)=r$ on the other hand.
@@
  • Imperceptible: Prompt injection attacks are often detected easily by inserting garbled code that can disrupt large models. We try to ensure that the adversarial text fits better into the problem context so that the model’s active defence mechanisms make it difficult to detect the presence of our prompt injections. The semantic constraint $\mathcal{D}(t^{\prime},t)$ can be replaced by a more relaxed constraint
    $\mathcal{D}(t^{\prime},t)<\epsilon.$ (2)
    That is, the distance $\mathcal{D}(\cdot,\cdot)$ between the clean text $t$ and the adversarial text $t^{\prime}$ is less than a given small threshold $\epsilon$ with $\epsilon>0$.
@@
  • Input-dependent: Compared to fixed triggers, input-dependent triggers are imperceptible and difficult to detect from humans in most cases [30]. It is obvious the adversarial text (or trigger) $t^{\prime}$ is input-dependent by Eqn. (2). We will insert trigger $t^{\prime}$ into $t$ through a prompt injection manner to form an attack prompt.
3.2 Analysis on Objective
@@
Below, we first discuss the necessary conditions for the LLM model to output different values ($M(t)\neq M(t^{\prime})$) under the conditions of clean text $t$ and adversarial text $t^{\prime}$, respectively.
@@
As can be seen from the problem (1), the input $t$ and output $r$ of model $M$ are both texts. To facilitate analysis, we discuss the embedded representation of texts in the embedding space. Assume that two different text inputs, $t$ and $t^{\prime}$, are represented by vectors $x=w(t)$ and $x^{\prime}=w(t^{\prime})$, respectively. The corresponding outputs $r$ and $r^{\prime}$ are represented by vectors $y=w(r)$ and $y^{\prime}=w(r^{\prime})$, where $w(\cdot)$ is a bijective embedded function from text to vector. Now, we re-formulate the output of the LLM model $M$ as a maximization problem of posterior probability in an enumerable discrete space $\mathcal{Y}$
$y=\operatorname*{arg\,max}_{\hat{y}\in\mathcal{Y}}p(\hat{y}|x)=w(M(w^{-1}(x))),y^{\prime}=\operatorname*{arg\,max}_{\hat{y}\in\mathcal{Y}}p(\hat{y}|x^{\prime})=w(M(w^{-1}(x^{\prime}))),$ (3)
where $w^{-1}(\cdot)$ is the inverse function of $w(\cdot)$ function. Obviously, we have
@@
That is, LLM has different posterior probability distributions under different input conditions, which is a necessary condition for LLM to output different values such that $M(t)\neq M(t^{\prime})$. To increase the likelihood that LLM will output different values, we quantify the divergence between the probability distributions $p(y|x)$ and $p(y|x^{\prime})$ with Kullback-Leibler (KL) divergence and maximize it
@@
Figure 1: Assumption that the output $y$ of LLM under the condition of $x$ satisfies the discrete Gaussian distribution: Answers (output) $y$ that are close to question (input) $x$ are usually more relevant to $x$ and have a higher probability of being sampled.
@@
as shown in Fig. 1, that is, the output of LLM is defined on a limited candidate set $\mathcal{Y}$, although this candidate set $\mathcal{Y}$ may be very large. Since the input $x$ and output $y$ of LLM are questions and answers, the spaces where they are located generally do not intersect with each other, so there are $y=\operatorname*{arg\,max}_{\hat{y}\in\mathcal{Y}}p(\hat{y}|x)\neq x$, which is different from the commonly used continuous Gaussian distribution $y=x$, as can be seen in Fig. 1. For the same question $x$, LLM usually outputs different answers and the answer $y$ most relevant to $x$ has a higher probability of being sampled. In the embedding space, the distance between $y$ and $x$ is usually closer. Similarly, answers $y$ that are almost uncorrelated with $x$ and that are far away from $x$ in the embedding space have a smaller probability of being sampled.
In order to solve the problem (6) theoretically, we do the following processing: 1) Replace the discrete Gaussian distribution with a continuous Gaussian distribution to facilitate the calculation of KL divergence although the output text on condition of input text still obey discrete distribution in practical applications; (2) The attack vector $x^{\prime}$ is constrained as a unit vector to preclude engagement with trivial scenarios, where we also care more about the direction of text embedding vectors in NLP. Subsequently, we present the following theorem (see Appendix Section A.2 and A.1).
Theorem 3.1
Assuming that variable $x^{\prime}$ is a unit vector, and $p(y|x)$ and $p(y|x^{\prime})$ respectively follows the Gaussian distribution $\mathcal{N}_{1}(y;x,\Sigma)$ and $\mathcal{N}_{2}(y;x^{\prime},\Sigma)$, then the maximization problem of $KL(p(y|x),p(y|x^{\prime}))$ is equivalent to the following optimization problem
@@
3.3 Solution of Problem (8) and Cos Similarity

### v2

@@ Abstract
Current large language models (LLMs) provide a strong foundation for large-scale user-oriented natural language tasks. Numerous users can easily inject adversarial text or instructions through the user interface, thus causing LLM model security challenges. Although there is currently much research on prompt injection attacks, most of these black-box attacks use heuristic strategies. It is unclear how these heuristic strategies relate to the success rate of attacks and thus effectively improve model robustness. To solve this problem, we redefine the goal of the attack: to maximize the KL divergence between the conditional probabilities of the clean text and the adversarial text. Furthermore, we prove that maximizing the KL divergence is equivalent to maximizing the Mahalanobis distance between the embedded representation $x$ and $x^{\prime}$ of the clean text and the adversarial text when the conditional probability is a Gaussian distribution and gives a quantitative relationship on $x$ and $x^{\prime}$. Then we designed a simple and effective goal-guided generative prompt injection strategy (G2PIA) to find an injection text that satisfies specific constraints to achieve the optimal attack effect approximately. It is noteworthy that our attack method is a query-free black-box attack method with a low computational cost. Experimental results on seven LLM models and four datasets show the effectiveness of our attack method.
Index Terms:
Prompt Injection, KL-divergence, LLM, Mahalanobis Distance.
I Introduction
@@
Large Language Models (LLMs) [1] are evolving rapidly in terms of architecture and applications. As they become more and more deeply integrated into our lives, the urgency of reviewing their security properties increases. A large number of previous studies [2] have shown that LLMs whose instructions are adjusted through reinforcement learning with human feedback (RLHF) are highly vulnerable to adversarial attacks. Therefore, studying adversarial attacks on large language models is of great significance, which can help researchers understand the security and robustness of large language models [3] and thus design more powerful and robust models to prevent such attacks.
@@
Attacks against large models are mainly divided into white-box attacks and black-box attacks. White-box attacks assume that the attacker has full access to the model weights, architecture, and training process to obtain gradient information, so the main method is gradient-based attacks. Guo et al. [4] proposed a new gradient-based distributed attack (GBDA), which uses the Gumbel-Softmax approximation technique to make the adversarial loss optimization differentiable on the one hand, and uses BERTScore and perplexity to enhance perceptibility and fluency on the other hand. Other white-box methods like DeepFool [5] and TextBugger [6] rely on calculating the model gradient to attack the model. TextBugger uses the gradient information to find the most effective text perturbation to complete the attack. DeepFool determines the minimum perturbation to find the boundary of sample classification error to achieve the attack goal. The disadvantage of these methods is that they can only attack large open-source language models. Still, for more widely used closed-source LLMs such as ChatGPT, these methods are powerless because they cannot obtain the architecture and parameters of the model.
@@
Compared with white-box attacks, black-box attacks limit attackers to access only API-type services. According to the attack granularity, black-box attacks [7] can be divided into letter level, word level, sentence level, and multi-level. Most black-box attacks use word replacement [8]: just find the most critical words in the text and replace them, or use some simple text enhancement methods [9], such as synonym replacement, random insertion, random exchange, or random deletion. Since black-box strategies cannot know the internal structure of large models, most attack methods use heuristic strategies. However, it is not clear how these heuristic strategies are related to the success rate of attacks, and how to effectively improve the success rate of attacks.
@@ II Related Work
In this section, we mainly review white-box and black-box attacks. Gradient-based attacks are in the white-box category, while token manipulation and prompt injection are all black-box attacks.
II-A Gradient-based Attack
@@
Gradient-based distribution attack (GBDA) [4] used Gumbel-Softmax approximation to make adversarial loss optimization differentiable on the one hand, and also uses BERTScore and perplexity to enhance perceptibility and fluency on the other hand. HotFlip [10] mapped text operations to vector space and calculated the derivative of the loss based on these vectors. Based on different tasks, AutoPrompt [11] used a gradient-based search strategy to find the most suitable prompt template for them. Autoregressive random coordinate ascent (ARCA) [12] searched for input-output pairs that match specific behavioral patterns under the broader optimization problem framework. Wallace et al. [13] proposed a gradient-guided token search method for searching for 5 tokens for short sequences, 1 for classification and 4 for generation, called universal adversarial triggers (UAT), which are used to trigger the model to produce specific predictions. However, gradient-based white-box attacks are not suitable for widely used closed-source large language models.
II-B Token Manipulation Attack
@@
Given a text input containing a sequence of tokens, simple word manipulations (e.g. replacing them with synonyms) can trigger the model to make incorrect predictions. Ribeiro et al. [14] manually defined adversarial rules (SEAR) that are semantically equivalent to heuristics to perform minimal token manipulations, thereby improving the attack success rate of the model. In contrast, Easy Data Augmentation (EDA) [15] defined a set of simpler and more general operations to augment the text, such as synonym replacement, random insertion, random swapping, or random deletion. BERT-Attack [16] replaces words with semantically similar ones through BERT based on the fact that context-aware prediction is a natural use case for masked language models. Unlike the above attack methods, our attack method attempts to insert generated adversarial prompts instead of just modifying certain words in the prompt.
II-C Prompt Injection Attack
@@
Generally, larger LLMs have excellent instruction tracking capabilities and are more vulnerable to attacks by malicious actors. An attacker [17] can easily embed instructions in the data and trick the model into misinterpreting it. In the work of Perez&Ribeiro [18], the goals of prompt injection attacks are divided into target hijacking and prompt leakage. The former attempts to redirect the original target of the LLM to a new target desired by the attacker, while the latter attempts to persuade the LLM to disclose the initial system prompt of the model program. Obviously, system prompts are very valuable to companies because they can significantly affect model behavior and change user experience. Liu et al. [19] found that LLMs are highly sensitive to escape characters and delimiters appearing in prompts, which implicitly convey instructions to start a new scope within the prompt by escaping. It is worth noting that our method for generating prompt injection attacks does not attempt to insert manually specified attack instructions, but instead attempts to influence the output of LLMs by generating confusing prompts based on the original prompts.
II-D Black-box Attack Paradigm
@@
The black box attack paradigm refers to a broad range of attack methods that exploit vulnerabilities in machine learning models. Boundary query-free attack BadNets [20] and boundary query-related attack [21] are two common black box attack methods, whose difference lies in whether the boundaries of the model need to be input in advance to observe the output of the model. Other black box attack methods include methods based on model substitution, such as practical black box attack[22] and context injection-based attacks, BERT-Attack [16].
III Methodology
III-A Threat Model with Black-box Attack
III-A1 Adversarial scope and goal.
@@ III-A Threat Model with Black-box Attack
Given a text $t$ containing multiple sentences, we generate a text $t^{\prime}$ to attack the language model, ensuring that the meaning of the original text $t$ is preserved, otherwise, then we believe that the attack text $t^{\prime}$ is attacking another text that is unrelated to $t$. Here, we use $\mathcal{D}(t^{\prime},t)$ to represent the distance between the semantics of text $t$ and $t^{\prime}$. If the LLM outputs $M(t)$ and $M(t^{\prime})$ differ, then $t^{\prime}$ is identified as an adversarial input for $M$. Our objective is formulated as follows:
$M(t)=r,\quad M\left(t^{\prime}\right)=r^{\prime},\quad\mathcal{D}(r,r^{\prime})\geq\varepsilon,\quad\mathcal{D}\left(t^{\prime},t\right)<\varepsilon,$ (1)
where the texts $r$ and $r^{\prime}$ are the outputs of model $M$ on text $t$ and $t^{\prime}$ respectively, and $r$ is also the groundtruth of text $t$. We introduce a distance function $\mathcal{D}(\cdot,\cdot)$ and a small threshold $\varepsilon$ to measure the semantic relationship between two texts. In our problem, we aim to provide the following attack characteristics
@@
  • Effective: Problem 1 shows that the attack model ensures a high attack success rate (ASR) with $\mathcal{D}(M(t^{\prime}),r)\geq\varepsilon$ on one hand and maintains high benign accuracy with $\mathcal{D}(M(t),r)<\varepsilon$ on the other hand.
@@
  • Imperceptible: Prompt injection attacks are often detected easily by inserting garbled code that can disrupt large models. We try to ensure that the adversarial text fits better into the problem context so that the model’s active defense mechanisms make it difficult to detect the presence of our prompt injections.
@@
  • Input-dependent: Compared to fixed triggers, input-dependent triggers are imperceptible and difficult to detect from humans in most cases [23]. It is obvious the adversarial text (or trigger) $t^{\prime}$ is input-dependent by Eqn. (1). We will insert trigger $t^{\prime}$ into $t$ through a prompt injection to form an attack prompt (see Sec. III-D).
III-B Analysis on Objective
@@
For the convenience of discussion, we regard the text generation of LLM as a classification problem, where the output will be selected from thousands of texts and each text is regarded as a category. Below, we first discuss the necessary conditions for the LLM model to output different values ($M(t)\neq M(t^{\prime})$) under the conditions of clean text $t$ and adversarial text $t^{\prime}$, respectively.
@@
As can be seen from the problem (1), the input $t$ and output $r$ of model $M$ are both texts. To facilitate analysis, we discuss the embedded representation of texts in the embedding space. Assume that two different text inputs, $t$ and $t^{\prime}$, are represented by vectors $x=w(t)$ and $x^{\prime}=w(t^{\prime})$, respectively. The corresponding outputs $r$ and $r^{\prime}$ are represented by vectors $y=w(r)$ and $y^{\prime}=w(r^{\prime})$, where $w(\cdot)$ is a bijective function between text and vector. Since we assume that the outputs $r$ and $r^{\prime}$ of LLM come from an enumerable discrete space and there is a one-to-one correspondence between text representations and vector representations, their vector representations $y$ and $y^{\prime}$ are also enumerable. Now, we re-formulate the output of the LLM model $M$ as a maximization problem of posterior probability in an enumerable discrete space $\mathcal{Y}$
$\begin{split}y&=\operatorname*{arg\,max}_{\hat{y}\in\mathcal{Y}}p(\hat{y}|x)=w(M(w^{-1}(x))),\\ y^{\prime}&=\operatorname*{arg\,max}_{\hat{y}\in\mathcal{Y}}p(\hat{y}|x^{\prime})=w(M(w^{-1}(x^{\prime}))),\end{split}$ (2)
where $w^{-1}(\cdot)$ is the inverse function of $w(\cdot)$ the function. Obviously, we have
@@
That is, LLM has different posterior probability distributions under different input conditions, which is necessary for LLM to output different values such as $M(t)\neq M(t^{\prime})$. To increase the likelihood that LLM will output different values, we quantify the divergence between the probability distributions $p(y|x)$ and $p(y|x^{\prime})$ with Kullback-Leibler (KL) divergence and maximize it
@@
Figure 1: Assumption that the output $y$ of LLM under the condition of $x$ satisfies the discrete Gaussian distribution: Answers (output) $y$ close to question (input) $x$ are usually more relevant to $x$ and have a higher probability of being sampled.
@@
as shown in Fig. 1, that is, the output of LLM is defined on a limited candidate set $\mathcal{Y}$, although this candidate set $\mathcal{Y}$ may be very large. Since the input $x$ and output $y$ of LLM are questions and answers, the spaces where they are located generally do not intersect with each other, so there are $y=\operatorname*{arg\,max}_{\hat{y}\in\mathcal{Y}}p(\hat{y}|x)\neq x$, which is different from the commonly used continuous Gaussian distribution $y=x$, as can be seen in Fig. 1. For the same question $x$, LLM usually outputs different answers and the answer $y$ most relevant to $x$ has a higher probability of being sampled. In the embedding space, the distance between $y$ and $x$ is usually closer. Similarly, answers $y$ that are almost uncorrelated with $x$ and far away from $x$ in the embedding space have a smaller probability of being sampled.
To solve the problem (5) theoretically, we do the following processing: Replace the discrete Gaussian distribution with a continuous Gaussian distribution to facilitate the calculation of KL divergence although the output text on condition of input text still obeys discrete distribution in practical applications. Subsequently, we present the following theorem (see Appendix Section A and B for details).
Theorem 1 Assuming $p(y|x)$ and $p(y|x^{\prime})$ respectively follows the Gaussian distribution $\mathcal{N}_{1}(y;x,\Sigma)$ and $\mathcal{N}_{2}(y;x^{\prime},\Sigma)$, then the maximization $KL(p(y|x),p(y|x^{\prime}))$ is equivalent to maximizing the Mahalanobis distance $(x^{\prime}-x)^{T}\Sigma^{-1}(x^{\prime}-x)$, which is further transformed into the following minimization optimization problem given the clean input $x$
@@
III-C Solving Problem (7) Approximately via Cos Similarity